Suppose that leather is sold in a perfectly competitive industry. The industry short-run supply curve (marginal cost curve) is P = MC = 3Q. The demand for leather hides is given by Q = 60 − P.

a. Find the equilibrium market price and quantity.

b. Suppose that the leather tanning releases bad stuff into waterways. The external marginal cost is $5 per unit. Calculate the socially optimal level of output and price for the tanning industry.

c. What are the total social benefit and total social cost at the market equilibrium?

d. What is the market surplus at the market equilibrium?

e. What are the total social benefit and total social cost at the social optimal point?

f. What is the market surplus at the social optimal point?

g. What is the deadweight loss?

Please use the graph to explain your answer.
Answer #1

Q = 60 - P, so P = 60 - Q

(a)

Setting P = MC,

60 - Q = 3Q

4Q = 60

Q = 15

P = 60 - 15 = 45

(b)

MSC = MC + MEC = 3Q + 5

Setting P = MSC,

60 - Q = 3Q + 5

4Q = 55

Q = 13.75

P = 60 - 13.75 = 46.25

(c)

When Q = 15,

Total social benefit = Total social cost = P x Q = 45 x 15 = 675

(d)

When Q = 0, P = 60

Consumer surplus (CS) = (1/2) x (60 - 45) x 15 = 7.5 x 15 = 112.5

When Q = 0, MC = 0

Producer surplus (PS) = (1/2) x (45 - 0) x 15 = 7.5 x 45 = 337.5

Market surplus = CS + PS = 112.5 + 337.5 = 450

[GRAPH]

In following graph, market equilibrium is at point A where Demand (P) intersects MC with price P0 and quantity Q0. Efficient outcome is at point B where Demand (P) intersects MSC with price P1 and quantity Q1. Deadweight loss is area ABC.

NOTE: As HOMEWORKLIB Answering Policy, 1st 4 parts have been answered.